2x+3y+z=7
1x-4y-z=-3
3x+1y-2z=14

Answer by checkley79(3341)

#1 2x+3y+z=7
#2 1x-4y-z=-3
#3 3x+1y-2z=14 add.
------------------
6x-2z=18
3x-z=9
-z=9-3x
z=3x-9
#1 2x+3y+3x-9=7
5x+3y=7+9
5x+3y=16 multiply by 2
#2 x-4y-(3x-9)=-3
-2x-4y+9=-3
-2x-4y=-3-9
2x+4y=12 multiply by -5 & add to #1.
10x+6y=32
-10x-20y=-60
--------------------
-14y=-28
y=-28/-14
y=2 ans.
#1 5x+3*2=16
5x+6=16
5x=16-6
5x=10
x=10/5
x=2 ans.
#3 3*2+2-2z=14
6+2-2z=14
8-2z=14
-2z=14-8
-2z=6
z=6/-2
z=-3 ans.
Proof:
# 2 2-4*2-(-3)=-3
2-8+3=-3
5-8=-3
-3=-3